OpenStudy (anonymous):

Which number is not in scientific notation? 1.1 ⋅ 10213 7.8 ⋅ 108 3.02 ⋅ 10−10 35 ⋅ 104

OpenStudy (anonymous):

@Michele_Laino can you please help me with this ?

OpenStudy (michele_laino):

hint: a number written using scientific notation has only one unit figure, for example, the subsequent number is written using scientific notation: \[\large 6.62 \cdot {10^{ - 27}}\]

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (anonymous):

idk im confused

OpenStudy (michele_laino):

int: the integer portion of a number, which is written using the scientific notion, has only one figure, which is the unit figure, now, the integer part of 35, for last number, is composed by two figures

OpenStudy (michele_laino):

hint*

OpenStudy (michele_laino):

1.1 is an integer number with 1 integer figure 7.8 is an integer number with 1 integer figure 3.02 is an integer number with 1 integer figure 35 is an integer number with 2 figures

OpenStudy (anonymous):

sorry im back

OpenStudy (anonymous):

i still dont understand this

OpenStudy (anonymous):

wait so the one with 2 figures is not scientific notation

OpenStudy (michele_laino):

a number is written in scientific notation, when it is like this: \[\Large 6.62 \cdot {10^{ - 27}}\] now, among the numbers which you have provided, what is the one which is not written in that form?

OpenStudy (anonymous):

D the last one

OpenStudy (michele_laino):

a number written using scientific notation has only one figure at the left of decimal point

OpenStudy (michele_laino):

so, correct!

OpenStudy (anonymous):

gosh pre- algebra is difficut to me.

OpenStudy (anonymous):

well kind of

OpenStudy (michele_laino):

for example, the subsequent number: \[\Large 258 \cdot {10^{ - 28}}\] is written using scientific notation? @Marc1313